# NV-Embed: Improved Techniques for Training LLMs as Generalist Embedding Models

### v1

@@ Abstract
Decoder-only large language model (LLM)-based embedding models are beginning to outperform BERT or T5-based embedding models in general-purpose text embedding tasks, including dense vector-based retrieval. In this work, we introduce the NV-Embed model with a variety of architectural designs and training procedures to significantly enhance the performance of LLM as a versatile embedding model, while maintaining its simplicity and reproducibility. For model architecture, we propose a latent attention layer to obtain pooled embeddings, which consistently improves retrieval and downstream task accuracy compared to mean pooling or using the last <EOS> token embedding from LLMs. To enhance representation learning, we remove the causal attention mask of LLMs during contrastive training. For model training, we introduce a two-stage contrastive instruction-tuning method. It first applies contrastive training with instructions on retrieval datasets, utilizing in-batch negatives and curated hard negative examples. At stage-2, it blends various non-retrieval datasets into instruction tuning, which not only enhances non-retrieval task accuracy but also improves retrieval performance. Combining these techniques, our NV-Embed model, using only publicly available data, has achieved a record-high score of 69.32, ranking No. 1 on the Massive Text Embedding Benchmark (MTEB) (as of May 24, 2024), with 56 tasks, encompassing retrieval, reranking, classification, clustering, and semantic textual similarity tasks. Notably, our model also attains the highest score of 59.36 on 15 retrieval tasks in the MTEB benchmark (also known as BEIR). We will open-source the model at: https://huggingface.co/nvidia/NV-Embed-v1.
@@ 1 Introduction
The embedding models built on bidirectional language models (Devlin et al., 2018; Raffel et al., 2020) have dominated the landscape for years (e.g., Reimers & Gurevych, 2019; Gao et al., 2021; Wang et al., 2022; Izacard et al., 2021; Ni et al., 2021), although one notable exception is Neelakantan et al. (2022). The most recent work by Wang et al. (2023b) demonstrates that decoder-only LLMs can outperform frontier bidirectional embedding models (Wang et al., 2022; Ni et al., 2021; Chen et al., 2023) in retrieval and general-purpose embedding tasks. However, previous leading efforts (Wang et al., 2023b; Meng et al., 2024) have depended on fine-tuning LLMs using large volumes of proprietary synthetic data from GPT-4, which is not readily available to the community.
@@
    For model training, we introduce a two-stage contrastive instruction-tuning method, starting with the pretrained Mistral-7B (Jiang et al., 2023). In the first stage, we apply contrastive training with instructions on retrieval datasets, utilizing in-batch negative and curated hard-negative examples. In the second stage, we blend carefully curated non-retrieval datasets into the stage-one training data. Since in-batch negative samples may be misleading for non-retrieval tasks, we disable in-batch negative training in stage two. This design not only improves the accuracy of classification, clustering, and semantic textual similarity tasks, but also surprisingly enhances retrieval performance. Note that our training data is entirely publicly available and does not include any synthetic data from proprietary models like GPT-4. Our model is also not fine-tuned from existing embedding models.  [^1] For example, SFR-Embedding is fine-tuned from E5-mistral-7b-instruct.
@@
    Combining all the techniques, our NV-Embed model sets a new record high score of 69.32 and ranks No. 1 (as of May 22, 2024) on the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2022) across 56 embedding tasks. It significantly outperforms the previous leading embedding models: E5-mistral-7b-instruct (score: 66.63) (Wang et al., 2023b), SFR-Embedding (score: 67.56) (Meng et al., 2024), and Voyage-large-2-instruct (score: 68.28) (Voyage-AI, 2024). Notably, our model also attained the highest score of 59.35 on 15 retrieval tasks within the MTEB, which is drawn from BEIR benchmark (Thakur et al., 2021).
We organize the rest of the paper in the following. In § 2, we discuss the related work. We present the architectural and training details in § 3.
@@ 2.2 Decoder-only LLM-based Embedding Models
Decoder-only LLMs (Brown et al., 2020) were believed to underperform bidirectional models on general-purpose embedding tasks because: i) unidirectional attention limits the representation learning capability, and ii) the scaling of LLMs leads to very high-dimension embeddings, which may suffer from the curse of dimensionality.
@@
Neelakantan et al. (2022) initializes the embedding models with pre-trained GPT-3 models (Brown et al., 2020) and applies continued contrastive training. The hidden state from the last layer corresponding to the special token <EOS> at the end of the sequence is taken as the embedding of the input sequence. The latest text-embedding-3-large obtains MTEB score 64.59 (OpenAI, 2024) Most recently, E5-Mistral (Wang et al., 2023b) (METB: 66.63) applies contrastive learning with task-specific instructions on Mistral 7B (Jiang et al., 2023). It begins to outperform the state-of-the-art bidirectional models on comprehensive embedding benchmarks (Muennighoff et al., 2022) by utilizing a massive amount of synthetic data from the proprietary GPT-4 model. LLM2Vec (BehnamGhader et al., 2024) (METB score: 65.01) tries to build the embedding model from LLMs while only using public available data, but it is still worse than E5-Mistral.
@@
Given the notable success of E5-Mistral, SFR-Embedding-Mistral (Meng et al., 2024) (METB: 67.56) further fine-tunes it on the blend of non-retrieval and retrieval datasets for improved accuracy on both tasks, which is closely related to our NV-Embed. However, there are the following key differences: 1) NV-Embed is trained from scratch on Mistral 7B LLM directly using public available data, and not dependent on other embedding model or proprietary synthetic data. Consequently, we introduce a new architecture that eliminates unnecessary causal attention mask and further improves the sequence pooling mechanism with latent attention layer. 2) SFR-Embedding-Mistral uses task-homogeneous batching, which constructs batches consisting exclusively of samples from a single task. In contrast, our NV-Embed uses well-blended batches consisting samples from all tasks to avoid potential “zigzag” gradient updates, which leads to a new record high score on both full MTEB and retrieval tasks compared to SFR-Embedding-Mistral.
There are other recent works. Gecko (Lee et al., 2024a) (METB: 66.31) attempts to distill a smaller bidirectional embedding model from a decoder-only LLM (Gemini et al., 2023) by generating synthetic paired data. It refines the data quality by retrieving a set of candidate passages for each query and relabeling the positive and hard negative passages using the LLM. In addition, GritLM (Muennighoff et al., 2024) (METB: 65.66) unifies text embedding and generation into a single model.
3 Method
@@ 4 Training Data
While recent embedding models (Wang et al., 2023b; Meng et al., 2024; Lee et al., 2024a) have utilized both public supervised datasets and proprietary synthetic data from GPT-4 (OpenAI, 2023) or Gemini (Gemini et al., 2023), we exclusively employ public datasets to demonstrate our model’s capability in embedding tasks. Our training procedure incorporates both retrieval and non-retrieval tasks, including classification, clustering, and semantic textual similarity datasets.
@@
$q^{+}_{\text{inst}}=\texttt{Instruct}:\texttt{task\_{definition}}\ \ \texttt{Query}:q^{+}$ (2)
The instruction templates for each task_definition are provided in Table 6 for training and Table 7 for evaluation. Note that we mask out the instruction tokens in the output embeddings during both training and evaluation, although they still impact the output due to self-attention. We do not add instruction prefixes to documents.
@@ 4.1 Public Retrieval Datasets
We adopt the retrieval datasets as follows: MS MARCO (Bajaj et al., 2016), HotpotQA (Yang et al., 2018), Natural Question (Kwiatkowski et al., 2019), PAQ (Lewis et al., 2021), Stackexchange (Stack-Exchange-Community, 2023), Natural language inference (Group et al., 2022), SQuAD (Rajpurkar et al., 2016), ArguAna (Wachsmuth et al., 2018), BioASQ (Tsatsaronis et al., 2015), FiQA (Maia et al., 2018), FEVER (Thorne et al., 2018). Typically, these datasets do not contain its own hardnegatives, necessitating the mining of such examples. To address this, we further finetune another encoder-based embedding model (Wang et al., 2022) to select the hardnegatives on those datasets. Refer to Table 6 for the number of samples used for training.
@@ 4.2 Public Non-retrieval Datasets
Besides retrieval datasets, we also utilize non-retrieval datasets from three sub-tasks in MTEB benchmark: classification, clustering and semantic similarity (STS). We pre-process these datasets to use the same format as retrieval datasets for contrastive training: instructed query $q^{+}_{\text{inst}}$ (containing query $q^{+}$), positive document $d^{+}$ and hard negative documents $d^{-}_{0},...,d^{-}_{n}$.
@@
We utilize the English training splits of various classification datasets from MTEB Huggingface datasets (Muennighoff et al., 2022; Lhoest et al., 2021). The classification datasets that we use are: AmazonReviews-Classification (McAuley & Leskovec, 2013), AmazonCounterfactual-Classification (O’Neill et al., 2021), Banking77-Classification (Casanueva et al., 2020), Emotion-Classification (Saravia et al., 2018), IMDB-Classification (Maas et al., 2011), MTOPIntent-Classification (Li et al., 2021), ToxicConversations-Classification (Adams et al., 2019), TweetSentimentExtraction-Classification (Maggie, 2020).
Because the training splits of Emotion-Classification and AmazonCounterfactual-Classification contain some content similar to their evaluation splits, we use BM25 (Robertson et al., 2009) similarity thresholds to remove similar content from the training splits before subsampling. We use the $text$ field as the $q^{+}$, $label\_text$ field as the $d^{+}$ and random sample among other $label\_text$ values for $d^{-}_{k}$. Since the AmazonReviewsClassification dataset does not provide the $label\_text$ field, we generate label texts associated with values in the $label$ field. For subsampling the classification datasets, we perform the stratified sampling across $d^{+}$.
We approach clustering in a similar manner as classification by employing the cluster labels for positives and negatives. We utilize the raw cluster label datasets raw_arxiv, raw_biorxiv and raw_medrxiv datasets from MTEB Huggingface datasets and filter out common content from the MTEB evaluation set of {Arxiv/Biorxiv/Medrxiv}-Clustering-{S2S/P2P} tasks. We use the $title$ field for $q^{+}$ for the S2S datasets and the $abstract$ field for $q^{+}$ for the P2P datasets. We use the $category$ field or random sample from the $categories$ field for $d^{+}$ and random sample other categories for $d^{-}_{k}$. We also use the raw label dataset for TwentyNewsgroups-Clustering (Lang, 1995) and remove any content that match with the MTEB evaluation set of the TwentyNewsgroups-Clustering task. For subsampling the clustering datasets, we perform stratified sampling across $d^{+}$.
@@
We use the training splits of three semantic similarity datasets STS12 (Agirre et al., 2012), STS22 (Chen et al., 2022), STS-Benchmark (Cer et al., 2017) from MTEB Huggingface datasets. For any pair of texts with associated relevance scores $(t_{a},t_{b},score)$, we create two examples $(q^{+}=t_{a},d^{+}=t_{b})$ and $(q^{+}=t_{b},d^{+}=t_{a})$ if $score\geq 4$. We mine the hard negatives $d^{-}_{k}$ from the pool of all texts using BM25, selecting the highest matching texts with rank >=2 that do not have relevance scores >2.5 with $q^{+}$.
Table 1: Top MTEB leaderboard models as of 2024-05-22. We use the original model names on the leaderboard for clarity.
Embedding Task Retrieval (15) Rerank (4) Cluter. (11) PairClass. (3) Class. (12) STS (10) Summ.( 1) Avg. (56)
Mertric nDCG@10 MAP V-Meas. AP Acc. Spear. Spear.
NV-Embed 59.36 60.59 52.80 86.91 87.35 82.84 31.2 69.32
NV-Embed (mean pool) 58.71 60.75 52.80 85.85 87.06 82.53 30.49 68.98
SFR-Embedding 59.00 60.64 51.67 88.54 78.33 85.05 31.16 67.56
Gte-Qwen1.5-7B-instruct 56.24 60.13 55.83 87.38 79.6 82.42 31.46 67.34
Voyage-lite-02-instruct 56.6 58.24 52.42 86.87 79.25 85.79 31.01 67.13
GritLM-7B 57.41 60.49 50.61 87.16 79.46 83.35 30.37 66.76
Google-gecko 55.7 58.9 47.48 87.61 81.17 85.07 32.63 66.31
LLM2Vec-Meta-Llama-3 56.63 59.69 46.45 87.79 75.92 83.58 30.94 65.01

### v2

@@ Abstract
Decoder-only large language model (LLM)-based embedding models are beginning to outperform BERT or T5-based embedding models in general-purpose text embedding tasks, including dense vector-based retrieval. In this work, we introduce the NV-Embed model, incorporating architectural designs, training procedures, and curated datasets to significantly enhance the performance of LLM as a versatile embedding model, while maintaining its simplicity and reproducibility. For model architecture, we propose a latent attention layer to obtain pooled embeddings, which consistently improves retrieval and downstream task accuracy compared to mean pooling or using the last <EOS> token embedding from LLMs. To enhance representation learning, we remove the causal attention mask of LLMs during contrastive training. For training algorithm, we introduce a two-stage contrastive instruction-tuning method. It first applies contrastive training with instructions on retrieval datasets, utilizing in-batch negatives and curated hard negative examples. At stage-2, it blends various non-retrieval into instruction tuning, which not only enhances non-retrieval task accuracy but also improves retrieval performance. For training data, we utilize the hard-negative mining, synthetic data generation and existing public available datasets to boost the performance of embedding model. By combining these techniques, our NV-Embed-v1 and NV-Embed-v2 models obtained the No.1 position on the Massive Text Embedding Benchmark (MTEB) (as of May 24, 2024 and August 30, 2024, respectively) across 56 embedding tasks, demonstrating the sustained effectiveness of the proposed methods over time. Additionally, it achieved the highest scores in the Long Doc section and the second-highest scores in the QA section of the AIR Benchmark, which covers a range of out-of-domain information retrieval topics beyond those in MTEB. We open-source the model at: https://huggingface.co/nvidia/NV-Embed-v2.
@@ 1 Introduction
The embedding models built on bidirectional language models (Devlin et al., 2018; Raffel et al., 2020) have dominated the landscape for years (e.g., Reimers & Gurevych, 2019; Gao et al., 2021; Wang et al., 2022; Izacard et al., 2021; Ni et al., 2021), although one notable exception is Neelakantan et al. (2022). The recent work by Wang et al. (2023b) demonstrates that decoder-only LLMs can outperform frontier bidirectional embedding models (Wang et al., 2022; Ni et al., 2021; Chen et al., 2023) in retrieval and general-purpose embedding tasks.
@@
    For model training, we introduce a two-stage contrastive instruction-tuning method, starting with the pretrained Mistral-7B (Jiang et al., 2023). In the first stage, we apply contrastive training with instructions on retrieval datasets, utilizing in-batch negative and curated hard-negative examples. In the second stage, we blend carefully curated non-retrieval datasets into the stage-one training data. Since in-batch negative samples are misleading for non-retrieval tasks in some cases, we disable in-batch negative training in stage two. This design not only improves the accuracy of classification, clustering, and semantic textual similarity tasks, but also surprisingly enhances retrieval performance. Note, our model is also not fine-tuned from existing embedding models [^1] For example, SFR-Embedding and Linq-Embed are fine-tuned from E5-mistral-7b-instruct..
@@
    Training data is one of the most crucial factors in achieving state-of-the-art results. We provide a detailed recipe on the curation of training datasets, including dataset-specific information, the positive-aware hard-negative mining technique to enhance contrastive training, the synthetic data generation and example-based multi-class labeling. This enables the community to easily reproduce and even surpass our model, ultimately advancing the development of the embedding models.
  4. 4.
    Our NV-Embed-v1 model obtained the No.1 position on the Massive Text Embedding Benchmark (MTEB) (as of May 24, 2024) (Muennighoff et al., 2022) across 56 embedding tasks. By improving the curation of the training data, NV-Embed-v2 model set a new record high score of 72.31 and reclaimed the No. 1 spot (as of Aug 30, 2024) on the highly competitive MTEB leaderboard, further demonstrating the sustained effectiveness of our approach. Note that our model also attains the highest scores in 15 retrieval tasks (commonly referred to as BEIR (Thakur et al., 2021)), 11 clustering tasks, and 12 classification tasks in the MTEB benchmark. See Table 1 for detailed information. Additionally, it secured the highest scores in Long Doc section and the second scores in QA section on the AIR-Benchmark which covers a range of out-of-domain information retrieval topics beyond those in MTEB.
We organize the rest of the paper in the following. In § 2, we discuss the related work. We present the architectural and training method in § 3. We provide detailed recipe of training data curation in § 4. We present the experiment results in § 5 and conclude the paper in § 6. AIR-bench results are shown in § A.
Table 1: Top MTEB leaderboard models as of ICLR submission date (2024-10-01). We use the original model names on the leaderboard for clarity.
Embedding Task Retrieval (15) Rerank (4) Cluster. (11) PairClass. (3) Class. (12) STS (10) Summ.( 1) Avg. (56)
Mertric nDCG@10 MAP V-Meas. AP Acc. Spear. Spear.
NV-Embed-v2 62.65 60.65 58.46 88.67 90.37 84.31 30.7 72.31
Bge-en-icl (zero shot) 61.67 59.66 57.51 86.93 88.62 83.74 30.75 71.24
Stella-1.5B-v5 61.01 61.21 57.69 88.07 87.63 84.51 31.49 71.19
SFR-Embedding-2R 60.18 60.14 56.17 88.07 89.05 81.26 30.71 70.31
Gte-Qwen2-7B-instruct 60.25 61.42 56.92 85.79 86.58 83.04 31.35 70.24
NV-Embed-v1 59.36 60.59 52.80 86.91 87.35 82.84 31.2 69.32
Bge-multilingual-gemma2 59.24 59.72 54.65 85.84 88.08 83.88 31.2 69.88
SFR-Embedding 59.00 60.64 51.67 88.54 78.33 85.05 31.16 67.56
GritLM-7B 57.41 60.49 50.61 87.16 79.46 83.35 30.37 66.76
@@ 2.2 Decoder-only LLM-based Embedding Models
Decoder-only LLMs (Brown et al., 2020) were believed to underperform bidirectional models on general-purpose embedding tasks for years, because: i) unidirectional attention limits the representation learning capability, and ii) the scaling of LLMs leads to very high-dimension embeddings, which may suffer from the curse of dimensionality.
@@
The early work by Neelakantan et al. (2022) initializes embedding models using pre-trained, decoder-only GPT-3 models (Brown et al., 2020) and applies continued contrastive training. The hidden state from the final layer, corresponding to the special token <EOS> at the end of the sequence, is used as the embedding for the input sequence. Its latest successor, text-embedding-3-large, achieves an MTEB score of 64.59 (OpenAI, 2024). Most recently, E5-Mistral (Wang et al., 2023b) (MTEB: 66.63) applies contrastive learning with task-specific instructions on Mistral 7B (Jiang et al., 2023). It begins to outperform the state-of-the-art bidirectional models on comprehensive embedding benchmarks (Muennighoff et al., 2022) by utilizing a massive amount of synthetic data from the proprietary GPT-4 model. LLM2Vec (BehnamGhader et al., 2024) (MTEB score: 65.01) tries to build the embedding model from LLMs while only using public available data, but it is still worse than E5-Mistral.
@@
Given the success of E5-Mistral, SFR-Embedding-Mistral (Meng et al., 2024b) (MTEB: 67.56) and SFR-Embedding-2R (Meng et al., 2024a) (MTEB: 70.31) further fine-tunes this model on the blend of non-retrieval and retrieval datasets for improved accuracy on both tasks, which is closely related to our NV-Embed. However, there are the following key differences: 1) NV-Embed is trained from scratch on Mistral 7B LLM directly using public available data, and not dependent on other embedding model or proprietary synthetic data. Consequently, we introduce a new architecture that eliminates unnecessary causal attention mask and further improves the sequence pooling mechanism with latent attention layer. 2) SFR-Embedding-Mistral uses task-homogeneous batching, which constructs batches consisting exclusively of samples from a single task. In contrast, our NV-Embed uses well-blended batches consisting samples from all tasks to avoid potential “zigzag” gradient updates, which leads to a new record high score on both full MTEB and retrieval tasks compared to SFR-Embedding-Mistral.
Over the past year, MTEB has become one of the most competitive leaderboards across all AI categories, leading to significantly increased competition among participants. Many of the recent top-performing models (e.g., stella-1.5B-v5, gte-Qwen2-7B-instruct, bge-multilingual-gemma2, voyage-large-2-instruct, and text-embed-3-large) have not disclosed key technical details necessary for reproduction, particularly the blend of training data used. Among the recently disclosed works, GritLM (Muennighoff et al., 2024) (MTEB: 65.66) unifies text embedding and generation into a single LLM model. In addition, bge-en-icl (Li et al., 2024) (MTEB: 71.24) enhances query embeddings by introducing few-shot examples on the query side, utilizing the in-context learning (ICL) capabilities in text embedding tasks. This approach introduces an overhead by supplying task-relevant examples to the query during the training process. To maintain zero-shot evaluation accuracy, both zero-shot and few-shot samples are included during training. In our paper, we focus on comparing the zero-shot evaluation accuracy of the bge-en-icl model to ensure the fair comparisons during the evaluation phase.
Another area of research focuses on improving data curation processes to enhance the accuracy of fine-tuning retrieval embedding models. Gecko (Lee et al., 2024a) (MTEB: 66.31) attempts to distill a smaller bidirectional embedding model from a decoder-only LLM (Gemini et al., 2023) by generating synthetic paired data. It refines the data quality by retrieving a set of candidate passages for each query and relabeling the positive and hard negative passages using the LLM. Linq-embed-mistral (Kim et al., 2024) utilized LLMs to refine data by generating, filtering, and mining negative samples. Meanwhile, NV-Retriever (Moreira et al., 2024) introduced a positive-aware hard-negative mining technique that considers positive relevance scores to more effectively eliminate false negatives. In this work, we apply this positive-aware hard-negative technique to curate the samples and enhance the contrastive training.
3 Methods
@@ 4 Training Data
For training data, we employ public retrieval and non-retrieval datasets and synthetically generated samples to demonstrate our model’s capability in embedding tasks. Our training procedure incorporates both retrieval and non-retrieval tasks including classification, clustering, and semantic textual similarity datasets.
@@
$q^{+}_{\text{inst}}=\texttt{Instruct}:\{\texttt{task\_{definition}}\}\ \ \texttt{Query}:q^{+}$ (2)
The instruction templates for each {task_definition} are provided in Table 9 for training and Table 10 for evaluation. Note, we mask out the instruction tokens in the output embeddings during both training and evaluation, although they still impact the output due to self-attention. We do not add any instruction prefix to document corpus.
@@ 4.1 Public Retrieval Datasets
We adopt the retrieval datasets as follows: MSMARCO (Bajaj et al., 2016), HotpotQA (Yang et al., 2018), Natural Question (Kwiatkowski et al., 2019), PAQ (Lewis et al., 2021), Stack Exchange (Stack-Exchange-Community, 2023), Natural Language Inference (Group et al., 2022), SQuAD (Rajpurkar et al., 2016), ArguAna (Wachsmuth et al., 2018), BioASQ (Tsatsaronis et al., 2015), FiQA (Maia et al., 2018), FEVER (Thorne et al., 2018), HoVer (Jiang et al., 2020), SciFact (Wadden et al., 2022), NFCorpus, MIRACL (Zhang et al., 2023) and Mr.TyDi (Zhang et al., 2021).
It is important to note that certain datasets (e.g., MSMARCO) are training splits of the MTEB Benchmark, which we follow the existing practices established by leading generalist embedding models (Meng et al., 2024b; Wang et al., 2023b; BehnamGhader et al., 2024; Muennighoff et al., 2024). Table 9 further provides the number of samples used for training. We demonstrate the zero-shot generalization capability of NV-Embed on AIR-bench in A.
4.1.1 Hardnegative mining technique
Embedding models are trained using contrastive learning (Gao et al., 2021), aiming to increase the similarity between the embeddings of a query and its relevant passages (positives) while reducing the similarity with irrelevant passages (negatives). Public retrieval datasets typically only contains the positive query-passage pairs but do not contain its own hardnegatives, making it necessary to mine of such negative examples. To address this, we apply the recently proposed positive-aware hard-negative technique (Moreira et al., 2024) that considers the positive relevance scores for better false negatives removal. Following the ablation studies in Moreira et al. (2024), we use E5-mistral-7b-instruct (Wang et al., 2023b) as a teacher retrieval model to identify the optimal hardnegative passages relevant to the query. We set the maximum threshold for negative scores based on a percentage of the positive score (TopKPercPos) with a 95% margin, described as follows: max_negative_score_threshold = pos_score * percentage_margin.
@@ 4.2 Public Non-Retrieval Datasets
Besides retrieval datasets, we utilize public non-retrieval datasets mainly from three sub-tasks in MTEB benchmark: classification, clustering and semantic similarity (STS). We pre-process the format of these datasets to become the compatible with retrieval datasets for contrastive training: query $q^{+}$, positive document $d^{+}$ and hard negative documents {$d^{-}_{0},...,d^{-}_{n}$}.
@@
For classification, we utilize the English training splits of various datasets from MTEB Huggingface datasets (Muennighoff et al., 2022; Lhoest et al., 2021). The classification datasets that we use are as follows: AmazonReviews (McAuley & Leskovec, 2013a), AmazonCounterfactual (O’Neill et al., 2021), Banking77 (Casanueva et al., 2020), Emotion (Saravia et al., 2018), IMDB (Maas et al., 2011), MTOPDomain/MTOPIntent (Li et al., 2021), ToxicConversations (Adams et al., 2019), TweetSentimentExtraction (Maggie, 2020), AmazonPolarity (McAuley & Leskovec, 2013b), MassiveScenario/MassiveIntent (FitzGerald et al., 2022). For the Emotion and AmazonCounterfactual classification datasets we use BM25 (Robertson et al., 2009) similarity thresholds to filter out training data that is similar to the MTEB evaluation set.
For clustering datasets, we utilize the raw_arxiv, raw_biorxiv and raw_medrxiv datasets from MTEB Huggingface datasets, TwentyNewsgroups (Lang, 1995), Reddit (Geigle et al., 2021), StackExchange (Geigle et al., 2021), RedditP2P (Reimers, 2021b) and StackExchangeP2P (Reimers, 2021a) We filter out any training data that match the MTEB evaluation set.
The classification and clustering datasets provide examples and corresponding class/cluster labels. The example texts extracted from the appropriate $text$/$title$/$abstract$ field are used for the query $q^{+}$. For binary classification tasks the label texts are used as documents $d^{+},d^{-}$. For multi-class classification and clustering tasks, a randomly sampled example from the ground-truth class/cluster is used for the positive document $d^{+}$ and randomly sampled examples from other classes/clusters are used for negative documents $d^{-}_{k}$. We will present ablation experiments supporting this approach in section 5.3.4.
@@
For semantic textual similarity datasets, we use the training splits of three semantic similarity datasets STS12 (Agirre et al., 2012), STS22 (Chen et al., 2022), STS-Benchmark (Cer et al., 2017) from MTEB Huggingface datasets. For any pair of texts with associated relevance scores $(t_{a},t_{b},score)$, we create two examples $(q^{+}=t_{a},d^{+}=t_{b})$ and $(q^{+}=t_{b},d^{+}=t_{a})$ if $score\geq 4$. We mine the hard negatives $d^{-}_{k}$ from the pool of other texts using the same technique as section 4.1.1. Task instructions are appended to $d^{+},d^{-}$ since they are symmmetric with the query.
4.3 Synthetic Tasks Dataset
Due to the limited variety of subjects and tasks in public training datasets, the available instruction templates for training are also restricted. To enhance task-wise generalization, we employ the Mixtral-8x22B-Instruct-v0.1 model (MistralAI, ) to create a dataset consisting of 120,000 synthetic examples across 60,000 synthetic tasks. Following a two-step prompting approach proposed by E5-mistral-7b-instruct (Wang et al., 2023b), we adjust the prompts for Mixtral-8x22B-Instruct-v0.1 and English text. We generate only the short-long, long-short, and short-short examples (40,000 of each), as we use public STS datasets and do not assess bitext retrieval tasks. Example prompts for synthetic data generation can be found in Appendix 12 and 13.